# DroneWiS: Automated Simulation Testing of small Unmanned Aerial Systems in Realistic Windy Conditions

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; October 27th–Nov 1st, 2024; California, USisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Application specific development environments
@@ 1. Introduction
Traditional sUAS simulation tools, like Gazebo and AirSim, use oversimplified wind models that ignore complex interactions with terrain and buildings. These models typically feature stochastic wind fluctuations or constant wind velocity, resulting in inaccurate simulations. Additionally, setting up realistic wind conditions is manual and labor-intensive, involving a 3D graphic designer to create the environment and a CFD engineer to compute wind flows using tools like OpenFoam. This process must be repeated for each simulation scenario, highlighting the need for automation to allow sUAS developers to create complex wind patterns more efficiently.
@@
The DroneWiS module uses state-of-the-art CFD software, OpenFOAM (Jasak, 2009), with a terrain scanning algorithm in Unreal Engine, to automatically simulate desired wind conditions in any given 3D environment. This innovative approach generates high-fidelity, dynamic wind vector data for various wind types, including uniform, turbulent, and multi-source winds.
@@ 2.1. Terrain Scanning
11:
12:  return  $voxelGrid$
13:  function ObstructionCheck($cellCenter$)
14:   Input: $cellCenter$
15:   Output: $bool$
16:   $faces\leftarrow$ Generate 6 faces of the voxel centered at $cellCenter$
17:  for each pair of opposite face centers in $faces$ do
18:      $hasLineOfSight\leftarrow$ lineTrace(face1, face2)
19:     if Not $hasLineOfSight$ then
20:
21:        return  true
22:     end if
23:  end for
24:
25:  return  false
26:  end function
@@ 2.2. Wind Flow Computation and Processing
After DroneWiS scans the environment, the resulting voxel grid is transferred to OpenFoam, a specialized CFD tool for processing voxel grids and computing wind vectors. DroneWiS automatically configures OpenFoam according to user-specified wind conditions, which include wind type (normal, turbulent, multi-source), wind direction, and wind velocity. Therefore, OpenFoam takes the voxel grid and wind characteristics as inputs to compute wind flows.
@@ 3. Application of DroneWiS in sUAS Testing
We recreated a 2022 FAA-reported incident where a drone crashed while taking off between two buildings (one taller than the other) to capture target imagery. Although the buildings shielded the takeoff location from the wind’s direct effects, as soon as the sUAS reached a higher altitude, high wind velocity blew the UAV off course significantly.
@@ 4. Conclusion
In this paper, we introduced DroneWiS, a novel component, integrated with the original DRV platform, that enables sUAS developers to conduct robust testing against complex wind patterns in a 3D simulation environment. By automating the computation and simulation of realistic wind conditions, DroneWiS provides a universally adaptable tool that can be integrated into various simulation platforms. Our comparative analysis of DroneWiS with AirSim’s wind simulations demonstrated significant improvements in wind effect fidelity, offering sUAS developers a powerful means to test and ensure safe operation of sUAS in real-world. In future work, we plan to adopt a machine learning-based approach to enhance the efficiency of wind vector generation for large-scale environments, support a broader range of wind conditions, and improve the fidelity of wind forces applied to sUAS.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3695351isbn: 979-8-4007-1248-7/24/10ccs: Software and its engineering Application specific development environments
@@ 1. Introduction
Traditional sUAS simulation tools, such as Gazebo and AirSim, use oversimplified wind models that ignore complex interactions with terrain and buildings. These models typically feature stochastic wind fluctuations or constant wind velocity, resulting in inaccurate simulations. Such oversights are particularly concerning because wind patterns change drastically near uneven terrain and buildings, posing significant safety risks to sUAS due to discrepancies between simulated and real-world wind flows. Additionally, computing wind flows using Computational Fluid Dynamics (CFD) in a 3D environment is a manual and labor-intensive process, involving a 3D scene designer to create the environment and a CFD engineer to compute wind flows using tools like OpenFoam (Jasak, 2009). This process must be repeated for each unique environment, highlighting the need for automation to allow sUAS developers to create complex wind patterns more efficiently.
@@
The DroneWiS module uses state-of-the-art CFD software, OpenFOAM (Jasak, 2009), with a terrain scanning algorithm in Unreal Engine, to automatically compute and simulate desired wind conditions in any given 3D environment. This allows sUAS developers to examine the effect of realistic wind flow around buildings and complex structures on sUAS trajectory. This innovative approach generates high-fidelity, dynamic wind vector data for various wind types, including uniform, turbulent, and multi-source winds.
@@ 2.1. Terrain Scanning
11:  return  $voxelGrid$
12:  function ObstructionCheck($cellCenter$)
13:   Input: $cellCenter$
14:   Output: $bool$
15:   $faces\leftarrow$ Generate 6 faces of the voxel centered at $cellCenter$
16:  for each pair of opposite face centers in $faces$ do
17:      $hasLineOfSight\leftarrow$ lineTrace(face1, face2)
18:     if Not $hasLineOfSight$ then
19:        return  true
20:     end if
21:  end for
22:  return  false
23:  end function
@@ 2.2. Wind Flow Computation and Processing
After DroneWiS scans the environment, the resulting voxel grid is transferred to OpenFoam for processing voxel grids and computing wind vectors. DroneWiS automatically configures OpenFoam according to user-specified wind conditions, which include wind type (normal, turbulent, multi-source), wind direction, and wind velocity. Therefore, OpenFoam takes the voxel grid and wind characteristics as inputs to compute wind flows.
@@ 3. Application of DroneWiS in sUAS Testing
We recreated a 2022 FAA-reported incident where a drone crashed while taking off between two buildings (one taller than the other) to capture target imagery. Although the buildings shielded the takeoff location from the wind’s direct effects, as soon as the sUAS reached a higher altitude, high wind velocity blew the UAV off course.
@@ 4. Conclusion
We introduced DroneWiS, a novel component, integrated with the original DRV platform, that enables sUAS developers to conduct simulation testing of sUAS against realistic, complex, and CFD-driven wind flows in a 3D environment. Our comparative analysis of DroneWiS with AirSim’s wind simulations demonstrated significant improvements in wind effect fidelity, offering sUAS developers a powerful means to test and ensure safe operation of sUAS in real-world. In future work, we plan to adopt a machine learning-based approach to enhance the efficiency of wind vector generation for large-scale environments, support a broader range of wind conditions, and improve the fidelity of wind forces applied to sUAS.